# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 4.4 Causal MPDAGs
Suppose $\mathcal{G}\in\mathfrak{G}$ satisfies $\mathcal{G}_{1}\preceq\mathcal{G}\preceq\mathcal{G}_{2}$ and differs from both. By the pseudo-rank condition, the skeletons of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are either identical or differ by one edge. If their skeletons coincide, then $\mathcal{G}_{2}$ has exactly one undirected edge that is directed in $\mathcal{G}_{1}$, so the intermediate $\mathcal{G}$ would need to both orient an undirected edge of $\mathcal{G}_{1}$ and undirect an edge of $\mathcal{G}_{2}$, which is impossible. If instead $\mathcal{G}$ has the skeleton of $\mathcal{G}_{1}$, which differs from that of $\mathcal{G}_{2}$ by one edge, then $\mathcal{G}_{2}$ contains one directed edge that does not exist in $\mathcal{G}_{1}$, while all other edges are shared; however, $\mathcal{G}$ must contain at least one undirected edge that is directed in $\mathcal{G}_{1}$ and can therefore not satisfy $\mathcal{G}\preceq\mathcal{G}_{2}$. The symmetric case, where $\mathcal{G}$ has the skeleton of $\mathcal{G}_{2}$ which differs from that of $\mathcal{G}_{1}$, follows analogously.
Proposition 12.
(MPDAG model complexity) For any MPDAG $\mathcal{G}\in\mathfrak{G}$ there exists a sequence $\mathcal{G}=x_{1}\gtrdot\dots\gtrdot x_{k}=\mathcal{G}_{\hat{0}}$ of MPDAGs, where $G_{\hat{0}}$ is the empty MPDAG and $k=\mathrm{pseudo}\text{-}\mathrm{rank}(\mathcal{G})$.
We prove Proposition 12 in Appendix G.2. Using this pseudo-rank function, we define $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ as the set of paths $x_{0}:=\mathcal{G}_{s},x_{1},\dots,x_{r}:=\mathcal{G}_{t}$ in the Hasse diagram corresponding to $\mathcal{L}$, such that each step satisfies $|\mathrm{pseudo}\text{-}\mathrm{rank}(x_{i})-\mathrm{pseudo}\text{-}\mathrm{rank}(x_{i-1})|=1$ for all $i$. In other words, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\subseteq P(\mathcal{G}_{s},\mathcal{G}_{t})$, where relative to $P(\mathcal{G}_{s},\mathcal{G}_{t})$, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ has the added constraint that the pseudo-rank changes by exactly one at each step in the path. This ensures that each transition $(x_{i-1},x_{i})$ between neighbor graphs corresponds to a minimal modification: either an undirected edge in one graph becomes directed in the other, or a directed edge in one graph is absent in the other. This constraint mimics graded posets, where neighbors differ by exactly one in rank, helping recover some useful structure in the non-graded setting of $\mathcal{L}$.
By Proposition 12, $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is a non-empty set. Since $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\subseteq P(\mathcal{G}_{s},\mathcal{G}_{t})$, we can therefore define a distance based on $P_{\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ that upper bounds $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ formalized as follows.
Corollary 4.
Let $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t}):=\min\{\allowbreak\mathrm{length}(p):p\in L_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})\}$, i.e., the length of the shortest path between $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$, among paths that lie inside $P_{\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$. Then, we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Corollary 4 has practical implications: the pseudo-distance $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is often significantly easier to compute than either the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ or the down-up and up-down distances $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$ and $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. In particular, recall that computing the model-oriented distance or the down-up and up-down distances requires enumerating neighbor graphs. However, as demonstrated in the example in Fig.˜6, the structure in neighbor graphs can vary widely, and no simple graphical characterization of neighborhood is known. In contrast, computing $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ only requires considering neighbors where the pseudo-rank changes by exactly one. This makes the enumeration process much simpler: for a graph $\mathcal{G}$, the relevant neighbors can be obtained by adding or removing any undirected edge, directing an undirected edge, or undirecting a directed edge — provided the resulting graph is a valid MPDAG in that it does not contain any induced subgraphs in Fig.˜2(b).
It may be tempting to claim $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is equal to $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$, the second variant of SHD between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. However, in the simple example among three nodes from earlier, $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is four, while $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=2$ is two. In fact, $\operatorname{\mathrm{SHD}}_{2}$ and consequently $\operatorname{\mathrm{SHD}}_{1}$ both lower bound $d_{\mathcal{L},\text{pseudo}}$ as formalized next.
Proposition 13.
For any pair of MPDAGs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, both variants of the SHD between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ is smaller than or equal to $d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{1},\mathcal{G}_{2})$; that is, $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{2})\leq d_{\mathcal{L},\text{pseudo}}(\mathcal{G}_{1},\mathcal{G}_{2})$.
We prove Proposition 13 in Appendix G.3. The bound in this proposition is used in the algorithm for computing $d_{\mathcal{L},\text{pseudo}}$; see Section 5.3. Note that while the SHD variants lower-bound $d_{\mathcal{L},\text{pseudo}}$, they do not always lower bound the model-oriented distance $d_{\mathcal{L}}$. For example, let $\mathcal{G}_{1}$ be the graph in Fig.˜6(a) and $\mathcal{G}_{3}$ be the graph in Fig.˜6(c). Then $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{3})=1$, while both SHD variants return n1n-1.
Figure 6: An MPDAG $\mathcal{G}_{1}$ in (a) and two of its neighboring MPDAGs in (b) and (c). The neighbor in (c) has n1n-1 more edges than $\mathcal{G}_{1}$.
4.4.1 Restriction to polytrees for MPDAGs
Similar to the case of CPDAGs, we can restrict our attention to a subclass of MPDAGs that are polytrees. Let $\mathfrak{G}_{\text{poly}}$ be the collection of MPDAGs over node set $V$ that are polytrees. Define the subposet $\mathcal{L}_{\text{poly}}=(\mathfrak{G}_{\text{poly}},\preceq)$ as the restriction of $\mathcal{L}$ to this subclass and let $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ denote the corresponding model-oriented distance. Let $d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$ be the analog of $d_{\mathcal{L},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$, but defined on the subposet $\mathcal{L}_{\text{poly}}$. Using a similar argument as Theorem 4, we can conclude that $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$. In fact, this inequality is tight: equality holds because the poset $\mathcal{L}_{\text{poly}}$ is graded with pseudo-rank as its ranking function. We formalize this next.
Theorem 4.
@@ 4.4.1 Restriction to polytrees for MPDAGs
The poset $\mathcal{L}_{\text{poly}}$ is graded with rank function $\mathrm{pseudo}\text{-}\mathrm{rank}(\cdot)$. Thus, the model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is equal to the pseudo-distance $d_{\mathcal{L}_{\text{poly}},\mathrm{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
We prove Theorem 4 in Appendix G.4. This result confirms that in the polytree setting, the pseudo-rank function defines a proper grading on $\mathcal{L}_{\text{poly}}$, and the corresponding distance exactly matches the model-oriented distance. Theorem 4 provides concrete justification for using the pseudo-rank-based distance more broadly.
Recall that when restricting CPDAGs to polytrees, the corresponding poset is lower semimodular, allowing us to appeal to Theorem 2 to conclude that the model-oriented distance equals the down-up distance. However, we show in Appendix E.3 that the poset obtained by restricting MPDAGs to polytrees is not lower semimodular. Thus, a similar simplification is not possible and we need to rely on the $A^{*}$-algorithm to compute $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ instead.
@@ 5 Algorithms for computing model-oriented distances
In this section, we develop algorithms for either computing or approximating the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for a pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$ from different graphical model classes $\mathfrak{G}$, including CPDAGs and MPDAGs. In Section 5.1, we present a general-purpose A* procedure for finding $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ across different graph classes. In Sections 5.2 and 5.3, we specialize this algorithm to the class of CPDAGs and MPDAGs, respectively, and present numerical experiments comparing the model-oriented distance to the two SHD variants $\operatorname{\mathrm{SHD}}_{1}$ and $\operatorname{\mathrm{SHD}}_{2}$ while tracking the computational cost of our algorithms.
5.1 General A* procedure
To compute the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ between two graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, we use an A* search algorithm, detailed in Algorithm 1. The search is performed on the poset $\mathcal{L}$. The algorithm maintains a priority queue (openSet) initialized with the start graph $\mathcal{G}_{s}$. This queue is ordered by a costs function $f(\mathcal{G})=g(\mathcal{G})+h(\mathcal{G})$, where: (i) $g(\mathcal{G})$ tracks the exact costs (i.e., number of steps) from $\mathcal{G}_{s}$ to $\mathcal{G}$, and (ii) $h(\mathcal{G})$ is a heuristic function that estimates the remaining cost from $\mathcal{G}$ to $\mathcal{G}_{t}$. To ensure that A* returns the distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$, the heuristic $h$ must be admissible:
Definition 9 (Admissible heuristic).
A heuristic $h(\cdot)$ is admissible if $h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for all $\mathcal{G}\in\mathfrak{G}$, i.e., it never overestimates the true distance.
We provide some more detail about Algorithm 1. At each iteration, the graph $\mathcal{G}_{\text{current}}$ with the smallest $f$-score is removed from the queue and expanded. If $\mathcal{G}_{\text{current}}=\mathcal{G}_{t}$, the algorithm terminates and returns $g(\mathcal{G}_{t})$. Otherwise, the algorithm examines all neighbors of $\mathcal{G}_{\text{current}}$ in $\mathcal{L}$. Here, the function $\mathrm{n}_{\mathcal{L}}(\cdot)$ finds neighboring graphs in the poset $\mathcal{L}$. For each neighbor $\mathcal{G}_{\text{neighbor}}$, it computes a tentative distance via $\mathcal{G}_{\text{current}}$. If this tentative distance is shorter than the path previously recorded to $\mathcal{G}_{\text{neighbor}}$, the algorithm updates $g(\mathcal{G}_{\text{neighbor}})$ and adds it to the queue with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$.
This A* framework is broadly applicable to any graph class $\mathfrak{G}$ that can be equipped with a connected poset $\mathcal{L}$. To use the algorithm in a given setting, one must provide: (i) an admissible heuristic that underestimates $d_{\mathcal{L}}(\mathcal{G}_{1},\mathcal{G}_{2})$ for all pairs $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathfrak{G}$, and (ii) a method for enumerating all neighbors of a given graph $\mathcal{G}\in\mathfrak{G}$, i.e., the graphs that cover or are covered by $\mathcal{G}$ in the poset.
input: Graph class $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, neighbor function $\mathrm{n}_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$
output : Model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$
1 begin
 $openSet\coloneqq\{\mathcal{G}_{s}\}$;
 // priority queue ordered by $g+h$
2 $g(\mathcal{G}_{s})\coloneqq 0$;

### v2

@@ 4.4 Causal MPDAGs
By Proposition˜10, given any $\mathcal{G}_{s},\mathcal{G}_{t}$, the set $\mathrm{P}_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ is non-empty because it contains the path that concatenates two maximal chains: first from $\mathcal{G}_{s}$ to $\mathcal{G}_{\hat{0}}$, then from $\mathcal{G}_{\hat{0}}$ to $\mathcal{G}_{t}$. The result then follows by definition. ∎
Due to the ease of enumerating graphs in the restricted neighborhood, $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ is easier to compute than $_{\mathcal{L}}, d,d_{\mathcal{L},\downarrow\uparrow}$ or $d_{\mathcal{L},\uparrow\downarrow}$. We also compare it with the structural Hamming distances in the next result.
Proposition 11 (MPDAG distance to SHD relationship).
For any MPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, it holds that $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{2})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{2})\leq d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{1},\mathcal{G}_{2})$.
We prove this result in Section˜G.2. See also Fig.˜S14 in the Appendix that compares these distances computed for polytree MPDAGs over 5 vertices. In Section˜5.3, we will use $\operatorname{\mathrm{SHD}}_{2}$ as a lower bound in the A* algorithm for computing $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$. Note that the structural Hamming distances do not lower bound the original $d_{\mathcal{L}}$. In the example of Fig.˜6, we have $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{1},\mathcal{G}_{3})=1$ but $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{1},\mathcal{G}_{3})=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{1},\mathcal{G}_{3})=n-1$.
4.4.1 Restriction to polytrees
We consider again restricting the class of graphs to $\mathfrak{G}_{\mathrm{poly}}$, which in this context denotes all MPDAGs that are polytrees over a vertex set $\mathrm{V}$. This induces the subposet $\mathcal{L}_{\mathrm{poly}}=(\mathfrak{G}_{\mathrm{poly}},\preceq)$ and further the corresponding shortest-path distance $d_{\mathcal{L}_{\mathrm{poly}}}$ between every pair of graphs in $\mathfrak{G}_{\mathrm{poly}}$. Furthermore, let $d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}$ be the analogue of $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ in Eq.˜7 but defined on the subposet $\mathcal{L}_{\mathrm{poly}}$ instead of $\mathcal{L}$. By an argument similar to the proof of Corollary˜4 (there still exists a path by concatenating two maximal chains of polytrees), we have $d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for every polytree $\mathcal{G}_{s},\mathcal{G}_{t}$. In fact, the result below shows that this inequality is an equality; see Section˜E.5 for its proof.
Theorem 7 (Polytree MPDAG poset).
Let $\mathcal{L}_{\mathrm{poly}}$ be the model-oriented poset associated with $(\mathfrak{G}_{\mathrm{poly}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}})$, where $\mathfrak{G}_{\mathrm{poly}}$ is the class of polytree MPDAGs over a finite vertex set $\mathrm{V}$. We have the following results.
  1. 1.
    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$ when $|\mathrm{V}|>2$.
  2. 2.
@@ 4.4.1 Restriction to polytrees
    $\mathcal{L}_{\mathrm{poly}}$ is graded with rank function equal to $\operatorname{\mathrm{pseudo-rank}}(\cdot)$.
  3. 3.
    $\mathcal{L}_{\mathrm{poly}}$ is lower semimodular.
  4. 4.
    It holds that
    $d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\mathrm{poly}}.$
  5. 5.
    It holds that
    $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})\leq d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t}),\quad\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}_{\mathrm{poly}}.$
@@ 5 Algorithms for computing model-oriented distances
In this section, we develop algorithms for computing the model-based distance. We will first present a generic A* algorithm suited to any graph class. Then, we will specialize the algorithm to probabilistic CPDAGs and causal, polytree MPDAGs in a manner that exploits their respective poset properties to save computation. Numerical results will be presented to illustrate computation time and to provide a comparison with the structural Hamming distance.
5.1 Generic A* algorithm with branch and bound
A* (Hart, Nilsson and Raphael, 1968) is a widely-used search algorithm to find the shortest path between two vertices on a graph. By employing a heuristic function to guide the search, it is typically more efficient than naive breadth-first or depth-first search. Suppose we want to find the distance between $\mathcal{G}_{s}$ (source) and $\mathcal{G}_{t}$ (target). Starting from $\mathcal{G}_{s}$, the algorithm maintains a priority queue (known as the open set) for the graphs to be visited, sorted by the index $f(\mathcal{G}):=g(\mathcal{G})+h(\mathcal{G})$ for each graph $\mathcal{G}$ in the queue. Here, $g(\mathcal{G})$ is the distance from $\mathcal{G}_{s}$ to $\mathcal{G}$ and $h(\mathcal{G})$ is the heuristic that estimates from below the distance from $\mathcal{G}$ to $\mathcal{G}_{t}$. For each iteration, a graph with the smallest index is popped from the queue and its neighbors are pushed to the queue, upon which their $f$ and $g$ values are updated if smaller values are found. The algorithm terminates when $\mathcal{G}_{t}$ is popped from the queue and returns $g(\mathcal{G}_{t})$ as the distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. When $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ are connected, A* is guaranteed to terminate. It can be shown that A* outputs the correct distance if the heuristic is admissible in the sense that $0\leq h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for every $\mathcal{G}\in\mathfrak{G}$. Depending on how the queue is implemented, A* reduces to breadth-first or depth-first search if a trivial heuristic $h\equiv 0$ is supplied. By using an admissible heuristic with smaller $d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})-h(\mathcal{G})$, A* becomes more efficient.
To further save computation with branch and bound, we also employ an upper bound $u(\mathcal{G})$ on the distance from $\mathcal{G}$ to $\mathcal{G}_{t}$ ($u\equiv+\infty$ is a trivial upper bound). By using $u(\cdot)$ to maintain $u^{\ast}$, the smallest upper bound on the distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ witnessed so far, the algorithm can terminate early whenever $f(\mathcal{G})\geq u^{\ast}$ due to the relation $f(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u^{\ast}$. Moreover, when expanding the neighborhood of $\mathcal{G}$, those graphs with $f$ no smaller than $u^{\ast}$ can be pruned immediately. Naturally, one can use $_{\mathcal{L},\uparrow\downarrow}, d,d_{\mathcal{L},\downarrow\uparrow}$ or their minimum as the upper bound. We present this generic A* in Algorithm˜1, which requires three subroutines: lower bound $h(\cdot)$, upper bound $u(\cdot)$ and $\textsc{EnumNeighbors}()$ that enumerates the neighborhood of a graph in the model-oriented poset. We establish the correctness of the algorithm in the next theorem, which we prove in Section˜I.1.
Theorem 8 (Algorithm correctness).
Suppose $\mathfrak{G}$ is a finite graph class and the associated model-oriented poset $\mathcal{L}=(\mathfrak{G},\preceq)$ is connected. For Algorithm˜1, suppose the subroutines $\textsc{EnumNeighbors}(\cdot)$, $u(\cdot)$ and $h(\cdot)$ satisfy the conditions specified at the beginning of the algorithm. Then, the algorithm will terminate and return $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Input: Graphs $\mathcal{G}_{s},\mathcal{G}_{t}$ in a finite graph class $\mathfrak{G}$ with model-oriented poset $\mathcal{L}=(\mathfrak{G},\preceq)$.
Require : Subroutine $\textnormal{{EnumNeighbors}}(\mathcal{G})$ that outputs the neighbors of $\mathcal{G}$ in $\mathcal{L}$;
subroutine $h(\mathcal{G})$ such that $0\leq h(\mathcal{G})\leq d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})$ for every $\mathcal{G}\in\mathfrak{G}$;
subroutine $u(\mathcal{G})$ such that $d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})\leq u(\mathcal{G})\leq+\infty$ for every $\mathcal{G}\in\mathfrak{G}$.
Output: Model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
1
[…]
// Upper bound on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$
$g\leftarrow\textnormal{{Dict}}[]$
// Length of the shortest path from $\mathcal{G}_{s}$ found so far
2 $g[\mathcal{G}_{s}]\leftarrow 0$
$\textnormal{{openSet}}\leftarrow\textnormal{{PriorityQueue}}\left((\mathcal{G}_{s},0),h(\mathcal{G}_{s})\right)$
// Item $(\mathcal{G}_{s},g[\mathcal{G}_{s}])$ with index $g[\mathcal{G}_{s}]+h(\mathcal{G}_{s})$